OpenStudy (czawlytko):

Help with hypothesis testing! A magazine is considering the launch of an online edition. The magazine will only go ahead with the launch if convinced that more than 25% of its current subscribers would subscribe. The magazine takes a simple random sample of 500 current subscribers and 137 express interest in the online version. Is there evidence that more than 25% of the magazine's current subscribers would subscribe to an online edition? Use a 5% significance level. Calculate the appropriate test-statistic for this test.

jimthompson5910 (jim_thompson5910):

You'll use this formula \[\Large z = \frac{\hat{p} - p}{SE}\] where \[\Large \text{Sample Proportion} =\hat{p} = \frac{137}{500} = 0.274\] \[\Large \text{Population Proportion} = p = 0.25\] \[\Large \text{Sample Size} = n = 500\]

jimthompson5910 (jim_thompson5910):

Also, \[\Large \text{Standard Error} = SE = \sqrt{\frac{p(1-p)}{n}}\] \[\Large SE = \sqrt{\frac{0.25(1-0.25)}{500}}\] \[\Large SE \approx 0.0193649\]

jimthompson5910 (jim_thompson5910):

I'll let you finish up. Let me know what you get or if you need more help.